OpenStudy (anonymous):

g(x) = 5x2 - 4x h(x) = -4x2 - 6x Find: (g - h)(3) How do I solve it with 2 different equations like this? I'm not asking for the answer, just need to know how to solve this one.

OpenStudy (anonymous):

step one find g(3) and h(3) step 2 (g-h)(3) = g(3) -h(3) Note to find g(3) replace x by 3 in equation g(x)

OpenStudy (anonymous):

So, would I put 3 where the x's are ?

OpenStudy (anonymous):

Then add the two equations together?

OpenStudy (anonymous):

in the g(x) g(x)=5(3)^2-4(3)

OpenStudy (anonymous):

g(3)=5(3)^2-4(3)

OpenStudy (anonymous):

g(3)=45-12=33

OpenStudy (anonymous):

Okay, I got it. So, you get the answer for those two, then subtract the two answers.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you . I did all of this last year, but I took homeschooling at the beginning of this year, and forgot how to do it all. But, thank you so much!

OpenStudy (anonymous):

welcome